OpenStudy (anonymous):

Simplify (Square root 3)/2 times (Square root 2)/2

OpenStudy (mertsj):

\[\frac{\sqrt{3}}{2}\times \frac{\sqrt{2}}{2}=\frac{\sqrt{6}}{4}\]

OpenStudy (anonymous):

thanks can you help me with one more?

OpenStudy (mertsj):

probably